An airplane starting from rest takes 25 s and 500 m of runway at constant acceleration to leave the ground solve its velocity when it becomes airborne.
1
Expert's answer
2020-08-04T16:00:01-0400

The distance traveled by an airplane

"d=\frac{v+v_0}{2}t"

Hence, the final velocity of an airplane

"v=\frac{2d}{t}-v_0=\frac{2\times 500\:\rm m}{25\:\rm s}-0\:\rm m/s=40\:\rm m/s"